A group of people were asked if they had run a red light in the last year. The number of people that responded yes is 354 and no is 281. Find the probability that if a person is chosen at random, he/she has run a red light in the last year. Round to two decimal places.

1
Expert's answer
2021-11-22T14:56:41-0500

"P(yes)=\frac{354}{354+281}=\frac{354}{635}=0.56."